# EarthquakeNPP: A Benchmark for Earthquake Forecasting with Neural Point Processes

### v2

@@ 4 Benchmarking Experiment
We use EarthquakeNPP to benchmark five spatio-temporal Neural Point Processes (NPPs), each with prior positive claims in earthquake forecasting, against the ETAS model described in Section 2.2.
Neural Spatio-Temporal Point Process (NSTPP) [5]: A probability density function (pdf)-based NPP that parametrizes the spatial pdf with continuous-time normalizing flows (CNFs). We evaluate their Attentive CNF model due to its superior computational efficiency and overall performance compared to the Jump CNF model [5].
Deep Spatio-Temporal Point Process (DeepSTPP) [97]: A conditional intensity function-based NPP that constructs a non-parametric space-time intensity function driven by a deep latent process. This model features a closed-form intensity function, eliminating the need for numerical approximations.
Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) [96]: A conditional intensity function-based NPP that jointly models the 3D space-time integral of the intensity and its derivative (the intensity function) using a dual network approach.
Spatio-temporal Diffusion Point Process (DSTPP) [94]: A generative NPP that does not have a likelihood function. DSTPP employs diffusion models to capture complex spatio-temporal dynamics.
Score Matching-based Pseudolikelihood Estimation of Neural Marked Spatio-Temporal Point Process (SMASH)[36]: A generative NPP that also lacks a likelihood function. SMASH adopts a normalization-free objective by estimating the pseudolikelihood of marked STPPs through score-matching.
@@ 4.1 Likelihood Evaluation
Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development [53, 23]. Other traditional next-event metrics like Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are misleading for earthquake forecasting [27], as earthquake occurrence follows power law distributions [33, 10] that are heavy-tailed, making the errors non-Gaussian and non-Laplacian — contrary to the assumptions underlying RMSE and MAE (see Section G).
@@
Unlike the NPPs, which follow the standard machine learning procedure of training, validation, and testing, ETAS does not typically incorporate validation in its estimation procedure. Thus, it is fit using both the training and validation windows combined. For NPPs, the likelihood for training, validation, and testing depends on events occurring before the respective splits through memory in their history. The exception is NSTPP, which lacks a direct dependency on prior events. Nevertheless, its likelihood is evaluated on the same data as the other models. The definition of the ETAS model (equation 4) specifies how the magnitudes of earthquakes in the history contribute towards the intensity function. This earthquake magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modeling choices beyond the scope of this work.
@@
The ETAS model consistently achieves the highest temporal log-likelihood, with NPPs performing comparably or, in some cases, marginally better, except at the higher magnitude thresholds of the SCEDC catalog. Among the NPPs, AutoSTPP and NSTPP perform well across several datasets, though their performance is more variable than that of DeepSTPP, which demonstrates consistent, comparable performance to ETAS. Closer examination of model performance over time (Figure 9) reveals that relative performance to ETAS is poorest during large earthquake sequences. This is likely due to ETAS leveraging the magnitude feature of the data, which enables it to handle these sequences effectively. Conversely, model performance is strongest during "background" periods, when no large earthquakes occur. During these periods, ETAS models the background with a constant rate, while the NPPs improve upon this by capturing the non-stationary nature of the background data. This effect is most pronounced in the ComCat dataset, which includes more complex physical processes, such as those near the Mendocino Triple Junction [24]. The improved relative temporal performance of all NPPs compared to ETAS, particularly when the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, indicates that low magnitude earthquakes provide valuable predictive information for NPPs.
@@
ETAS consistently outperforms all NPPs in spatial log-likelihood. As with temporal performance, relative performance to ETAS is weakest during large earthquake periods, likely due to the absence of a magnitude feature in the NPPs. AutoSTPP achieves the highest spatial log-likelihood, attributed to its ability to capture anisotropic Hawkes kernels (see Figure 2 of Zhou and Yu [96]), which are commonly observed in earthquake data [58].
Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Figure 3: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
@@ 4.2 CSEP Consistency Tests
EarthquakeNPP supports the earthquake forecast evaluation protocol developed by the Collaboratory for the Study of Earthquake Predictability (CSEP). In this procedure, a model generates 24-hour forecasts through 10,000 repeated simulations of earthquake sequences at the beginning of each day in the testing period. This approach mirrors how earthquake forecasts are produced in operational settings [86]. Models are then evaluated by comparing the observed sequence with the distribution of forecasts generated by the simulations. Four test statistics assess the temporal, spatial, likelihood, and magnitude components of the forecasts. A test is considered failed if the observed statistic falls within a pre-defined rejection region (Figure 11). We apply this evaluation procedure to the two generative NPPs (DSTPP and SMASH) alongside ETAS (Table 2) and present a case study on the 2010 M7.2 El Mayor-Cucapah earthquake, using the forecasts from these models (Figure 4). Appendix E provides an introduction to the CSEP consistency tests, with further details found at https://cseptesting.org/.
Table 2: CSEP consistency tests evaluate the calibration of daily forecasts from three models (ETAS, SMASH, DSTPP) on EarthquakeNPP datasets. A test is performed at the $\alpha=0.05$ significance level on each day in the testing period. The pass rate indicates the proportion of testing days with non-rejected hypotheses. If the model is the data generator, quantile scores should be uniformly distributed. The KS-Statistic quantifies deviation from uniformity (see Appendix E). ETAS is the only model that forecasts earthquake magnitudes, so is the only model evaluated with the magnitude test.
@@
ComCat ETAS 85.1% 0.222 92.0% 0.029 97.6% 0.128 93.8% 0.113
SMASH 72.4% 0.212 68.6% 0.217 87.6% 0.134
DSTPP 70.4% 0.116 88.6% 0.070 86.3% 0.138
SCEDC ETAS 81.7% 0.347 88.3% 0.119 95.9% 0.233 90.0% 0.256
SMASH 59.9% 0.602 51.1% 0.471 68.0% 0.611
DSTPP 0.0% 0.840 6.1% 0.935 0.8% 0.988
QTM_SanJac ETAS 88.8% 0.151 91.7% 0.095 96.6% 0.123 94.8% 0.076
SMASH 67.8% 0.290 55.6% 0.385 53.4% 0.584
DSTPP 78.4% 0.110 85.7% 0.240 85.3% 0.136
QTM_SaltonSea ETAS 77.8% 0.210 90.9% 0.206 96.4% 0.119 90.6% 0.136
SMASH 58.6% 0.486 53.6% 0.371 73.7% 0.451
DSTPP 69.6% 0.154 88.8% 0.136 85.6% 0.127
White ETAS 83.4% 0.167 86.6% 0.225 90.8% 0.233 88.8% 0.052
SMASH 61.3% 0.274 84.5% 0.150 67.7% 0.246
DSTPP 0.0% 0.987 30.9% 0.691 32.3% 0.892
@@ 5 Discussion and Conclusion
In benchmarking five NPP models against ETAS, we found that none outperformed the baseline, suggesting current NPPs are not yet suitable for operational use. While NPPs often perform well during low-activity background periods, they struggle during active phases following large earthquakes. This may be due to their lack of explicit magnitude dependence — a key feature in ETAS — and their limited memory, as NPPs encode only a fixed-length history. In contrast, ETAS sums over all past events, allowing long-past earthquakes to influence future rates. Future NPPs should incorporate magnitude information and develop mechanisms for longer-term memory.
@@
EarthquakeNPP, available at https://github.com/ss15859/EarthquakeNPP/tree/main, provides a platform for future NPP developments to be benchmarked against these initial results. The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions. Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
Acknowledgments and Disclosure of Funding
@@
This project is funded by Compass - Centre for Doctoral Training in Computational Statistics and Data Science (EPSRC Grant Ref EP/S023569/1). Compass is funded by United Kingdom Research and Innovation (UKRI) through the Engineering and Physical Sciences Research Council (EPSRC), https://www.ukri.org/councils/epsrc. This project also has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant 821115, Real-time earthquake rIsk reduction for a reSilient Europe (RISE), http://www.rise-eu.org) and by United States Geological Survey (USGS) EHP grant G24AP00059.
@@ A.1 Earthquake Catalog Generation
Data missingness, referred to in seismology as catalog (in)completeness, is the primary challenge faced with earthquake catalogs. It is an important and unavoidable feature, and is a result of how earthquakes are detected and characterised. Below, we briefly overview the process of generating an earthquake catalog to illustrate the data quality issues. In the subsequent section, we review catalog incompleteness and its potential impact on the performance and evaluation of forecast models.
@@
Seismometers and Seismic Networks. A seismometer is an instrument that detects and records the vibrations caused by seismic waves [75, 71]. It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities. Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., [92] and references therein). High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
@@
From Waveforms to Phase Picking. The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake. Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes [2]. Recent algorithms, such as machine learning classifiers [e.g. 98, 35] and template-matching [e.g. 62], can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
@@
Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake. Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event. Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms [82, 37]. Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist [e.g. 92, and references therein]. Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
@@
Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter [61]. The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers. This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons. This led to introduction of the moment magnitude scale (Mw) [22], which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
@@
where rigidity is a mechanical property of the rock along the fault, rupture area is the area of the fault that slipped, and slip is the distance the fault moved. Mw is determined seismologically via a spectral fitting process to the earthquake waveforms. In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales [e.g. 26, and references therein].

### v3

@@ 4 Benchmarking Experiment
We use EarthquakeNPP to benchmark five spatio-temporal Neural Point Processes (NPPs) against the ETAS model described in Section 2.2. Each of these NPPs has prior positive claims in earthquake forecasting, yet lacks performance comparison with the ETAS model.
Neural Spatio-Temporal Point Process (NSTPP) (Chen et al., 2021): A probability density function (pdf)-based NPP that parametrizes the spatial pdf with continuous-time normalizing flows (CNFs). We evaluate their Attentive CNF model due to its superior computational efficiency and overall performance compared to the Jump CNF model (Chen et al., 2021).
Deep Spatio-Temporal Point Process (DeepSTPP) (Zhou et al., 2022): A conditional intensity function-based NPP that constructs a non-parametric space-time intensity function driven by a deep latent process. This model features a closed-form intensity function, eliminating the need for numerical approximations.
Automatic Integration for Spatiotemporal Neural Point Processes (AutoSTPP) (Zhou and Yu, 2024): A conditional intensity function-based NPP that jointly models the 3D space-time integral of the intensity and its derivative (the intensity function) using a dual network approach.
Spatio-temporal Diffusion Point Process (DSTPP) (Yuan et al., 2023): A generative NPP that does not have a likelihood function. DSTPP employs diffusion models to capture complex spatio-temporal dynamics.
Score Matching-based Pseudolikelihood Estimation of Neural Marked Spatio-Temporal Point Process (SMASH)(Li et al., 2023): A generative NPP that also lacks a likelihood function. SMASH adopts a normalization-free objective by estimating the pseudolikelihood of marked STPPs through score-matching.
Appendix D provides details on the computational cost of training and inference for all the models tested.
@@ 4.1 Likelihood Evaluation
Since generating repeated sequences over forecast horizons is computationally costly, the seismology community uses the mean log-likelihood on held-out data for a more streamlined metric during model development (Ogata, 1988; Harte, 2015). Other traditional next-event metrics like Root Mean Square Error (RMSE) and Mean Absolute Error (MAE) are misleading for earthquake forecasting (Hodson, 2022), as earthquake occurrence follows power law distributions (Kagan, 1994; Felzer and Brodsky, 2006) that are heavy-tailed, making the errors non-Gaussian and non-Laplacian, contrary to the assumptions underlying RMSE and MAE (see Section I).
Figure 2: Test temporal log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. SCEDC_20, SCEDC_25 and SCEDC_30 correspond to magnitude thresholds (Mw 2.0, 2.5, 3.0) of the SCEDC dataset. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
Figure 3: Test spatial log-likelihood scores for all the spatio-temporal point process models on each of the EarthquakeNPP datasets. SCEDC_20, SCEDC_25 and SCEDC_30 correspond to magnitude thresholds (Mw 2.0, 2.5, 3.0) of the SCEDC dataset. Error bars of the mean and standard deviation are constructed for the NPPs using three repeat runs.
@@
Unlike the NPPs, which follow the standard machine learning procedure of training, validation, and testing, ETAS does not typically incorporate validation in its estimation procedure. Thus, it is fit using both the training and validation windows combined. For NPPs, the likelihood for training, validation, and testing depends on events occurring before the respective splits through memory in their history. The exception is NSTPP, which lacks a direct dependency on prior events. Nevertheless, its likelihood is evaluated on the same data as the other models.
To ensure that fitting ETAS on both the training and validation windows does not bias the comparison, we also tested an alternative configuration where ETAS was trained only on the training window. As shown in Appendix C, ETAS performance remains effectively unchanged under this setup. The ETAS formulation (Equation 4) also specifies how the magnitudes of prior earthquakes contribute to the conditional intensity; this magnitude dependence is not implemented in any of the NPPs we benchmark, since it requires modelling choices beyond the scope of this work.
The ETAS model consistently achieves the highest temporal log-likelihood, with NPPs performing comparably or, in some cases, marginally better, except at the higher magnitude thresholds of the SCEDC catalog. Among the NPPs, AutoSTPP and NSTPP perform well across several datasets, though their performance is more variable than that of DeepSTPP, which demonstrates consistent, comparable performance to ETAS. Differences in Poisson performance across Figures 2 and 3 are driven by variations in clustering strength, with weakly clustered catalogs appearing nearly Poisson-like and strongly clustered catalogs exhibiting larger departures.
@@
Closer analysis of model performance over time (see Section E) reveals that relative performance to ETAS is poorest during large earthquake sequences. This is likely due to ETAS leveraging the magnitude feature of the data, which enables it to handle these sequences effectively. Conversely, model performance is strongest during "background" periods, when no large earthquakes occur. During these periods, ETAS models the background with a constant rate, while the NPPs improve upon this by capturing the non-stationary nature of the background data. The improved relative temporal performance of all NPPs compared to ETAS, particularly when the magnitude threshold is lowered from $3.0$ to $2.0$ in the SCEDC dataset, indicates that low magnitude earthquakes provide valuable predictive information for NPPs.
@@
ETAS consistently outperforms all NPPs in spatial log-likelihood. Further analysis of model performance over space (see Section E) shows relative performance to ETAS is weakest in the most active and clustered areas (see Figures 12 and 13), likely due to the absence of a magnitude feature in the NPPs. However, NPPs tend to perform more competitively in regions characterised by spatially complex or diffuse seismicity. AutoSTPP achieves the highest spatial log-likelihood, attributed to its ability to capture anisotropic Hawkes kernels (see Figure 2 of Zhou and Yu (2024)), which are commonly observed in earthquake data (Page and van der Elst, 2022).
@@ 4.2 CSEP Consistency Tests
EarthquakeNPP supports the earthquake forecast evaluation protocol developed by the Collaboratory for the Study of Earthquake Predictability (CSEP). In this procedure, a model generates 24-hour forecasts through 10,000 repeated simulations of earthquake sequences at the beginning of each day in the testing period. This approach mirrors how earthquake forecasts are produced in operational settings (van der Elst et al., 2022). Models are then evaluated by comparing the observed sequence with the distribution of forecasts generated by the simulations. Four test statistics assess the temporal, spatial, likelihood, and magnitude components of the forecasts. A test is considered failed if the observed statistic falls within a pre-defined rejection region (Figure 14). We apply this evaluation procedure to the two generative NPPs (DSTPP and SMASH) alongside ETAS (Table 2) and present a case study on the 2010 M7.2 El Mayor-Cucapah earthquake, using the forecasts from these models (Figure 4). Appendix F provides an introduction to the CSEP consistency tests, with further details found at https://cseptesting.org/, and Appendix G provides further analysis on the simulated forecasts.
Table 2: CSEP consistency tests evaluate the calibration of daily forecasts from three models (ETAS, SMASH, DSTPP) on EarthquakeNPP datasets. A test is performed at the $\alpha=0.05$ significance level on each day in the testing period. The pass rate indicates the proportion of testing days with non-rejected hypotheses. If the model is the data generator, quantile scores should be uniformly distributed. The KS-Statistic quantifies deviation from uniformity (see Appendix F). ETAS is the only model that forecasts earthquake magnitudes, so is the only model evaluated with the magnitude test.
@@
ComCat ETAS 95.8% 0.222 92.0% 0.029 97.6% 0.128 93.8% 0.113
SMASH 86.2% 0.212 68.6% 0.217 87.6% 0.134
DSTPP 86.7% 0.116 88.6% 0.070 86.3% 0.138
SCEDC ETAS 92.6% 0.347 88.3% 0.119 95.9% 0.233 90.0% 0.256
SMASH 69.6% 0.602 51.1% 0.471 68.0% 0.611
DSTPP 27.4% 0.840 6.1% 0.935 0.8% 0.988
QTM_SanJac ETAS 95.4% 0.151 91.7% 0.095 96.6% 0.123 94.8% 0.076
SMASH 81.7% 0.290 55.6% 0.385 53.4% 0.584
DSTPP 85.7% 0.110 85.7% 0.240 85.3% 0.136
QTM_SaltonSea ETAS 93.6% 0.210 90.9% 0.206 96.4% 0.119 90.6% 0.136
SMASH 75.8% 0.486 53.6% 0.371 73.7% 0.451
DSTPP 87.7% 0.154 88.8% 0.136 85.6% 0.127
White ETAS 88.3% 0.167 86.6% 0.225 90.8% 0.233 88.8% 0.052
SMASH 69.3% 0.274 84.5% 0.150 67.7% 0.246
DSTPP 0.5% 0.987 30.9% 0.691 32.3% 0.892
@@
Further analysis of the simulated forecasts in Appendix G provides insight into the consistency test results. Temporally, all models struggle to capture the highest-rate seismicity days, indicating substantial room for improvement in modelling the most hazardous periods. SMASH exhibits highly variable, spiky daily rate forecasts that result in frequent over- and underprediction, while DSTPP produces much smoother forecasts that systematically underestimate seismicity across both background and active periods. Spatially, ETAS assigns concentrated rates along known fault structures through its explicit clustering mechanism. In contrast, SMASH generates diffuse spatial forecasts with weak contrast between active and inactive regions, whereas DSTPP more accurately follows fault-aligned structure but often assigns uniformly low spatial rates, particularly in the SCEDC and White datasets.
We were unable to apply the CSEP evaluation procedure for NSTPP, AutoSTPP and DeepSTPP, since the models are not explicitly formulated to be generative and therefore suffer from slow sampling (see details in Appendix D). This limitation significantly hinders their ability to be applied to real-time operational earthquake forecasting.
@@ 5 Discussion and Conclusion
In benchmarking five neural point process (NPP) models against ETAS, we found that none outperformed the baseline, indicating that current NPP architectures are not yet suitable for operational earthquake forecasting. While several NPPs achieve competitive performance during low-activity background periods, they consistently struggle during highly active phases following large earthquakes. Our results highlight several concrete architectural and methodological gaps, which we summarise below as actionable directions for future NPP development.
Action 1: Encode explicit magnitude dependence. ETAS explicitly encodes magnitude dependence, whereby larger earthquakes exponentially increase both the rate and spatial extent of subsequent seismicity. None of the benchmarked NPPs incorporate such explicit magnitude scaling, which limits their ability to capture the dominant influence of large events. Future NPP architectures could address this by introducing magnitude-aware design choices, such as hierarchical encodings that distinguish small and large events, magnitude-weighted attention mechanisms, or parameterisations aligned with the logarithmic frequency–magnitude scaling observed in seismicity (Richter, 1935). These approaches would allow NPPs to retain flexibility while incorporating structure that has proven critical for ETAS performance.
Action 2: Design scalable long-term memory mechanisms. All evaluated NPPs truncate the conditioning history due to the computational cost of sequence encoders, with models such as DeepSTPP and AutoSTPP conditioning on as few as 20 past events. In contrast, ETAS integrates the full event history, allowing long-past earthquakes, including large or spatially distant events, to influence future rates. Designing NPPs with scalable long-term memory is therefore a critical avenue for improvement. Promising directions include sparse or dilated attention mechanisms (Child et al., 2019; Hassani and Shi, 2022) to reduce quadratic complexity, hierarchical or coarse-to-fine representations of earthquake histories (Yang et al., 2016), and explicit memory compression modules (Kim et al., 2023) that preserve the influence of distant but significant events. Advances in long-context modelling within the NLP literature suggest that such mechanisms are technically feasible (Liu et al., 2025) and may translate naturally to earthquake triggering dynamics.
Action 3: Align generative training with operational evaluation. Third, our results reveal a mismatch between how generative NPPs are trained and how they are evaluated. Models such as SMASH and DSTPP are trained to predict or sample the next event, whereas CSEP consistency tests require simulating complete event sequences over fixed forecasting windows. This discrepancy helps explain why some generative models show reasonable short-term accuracy (Yuan et al., 2023; Li et al., 2023) but perform poorly in our multi-event simulation tests. Future generative NPPs may therefore benefit from training objectives that explicitly target long-horizon trajectory behaviour, for example by optimising multi-event simulation losses (e.g. Lüdke et al., 2024) or designing statistics (e.g. equation 13) aligned with the CSEP evaluations used in this study.
Action 4: Incorporate empirically supported scaling laws. Finally, our results suggest that the complete removal of physically motivated structure may be counterproductive. While NPPs aim to move beyond parametric models, ETAS kernels encode power-law scaling relationships that are strongly supported by empirical seismology (Kagan, 1994; Felzer and Brodsky, 2006). Hybrid architectures that combine neural density estimation with ETAS-inspired power-law kernels or magnitude-dependent triggering functions may offer a productive middle ground, retaining empirical laws while allowing greater flexibility than purely parametric formulations. For example, replacing Gaussian spatial kernels in existing NPPs (e.g in DeepSTPP) with learned power-law forms could improve their ability to represent aftershock clustering without sacrificing expressiveness.
@@
EarthquakeNPP, available at https://github.com/ss15859/EarthquakeNPP, provides a platform for future NPP developments to be benchmarked against these initial results. The platform is under ongoing development and in the future will see the direct comparison of emerging and other existing models developed within the seismology community, as well as an expansion of datasets included to other seismically active global regions. Successful NPP models on these datasets, for both log-likelihood and CSEP metrics, will be directly impactful to stakeholders in seismology, ultimately enabling their integration into operational earthquake forecasting by government agencies.
Acknowledgments
@@
This project is funded by Compass - Centre for Doctoral Training in Computational Statistics and Data Science (EPSRC Grant Ref EP/S023569/1). Compass is funded by United Kingdom Research and Innovation (UKRI) through the Engineering and Physical Sciences Research Council (EPSRC), https://www.ukri.org/councils/epsrc. This project also has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant 821115, Real-time earthquake rIsk reduction for a reSilient Europe (RISE), http://www.rise-eu.org) and by United States Geological Survey (USGS) EHP grants G24AP00059 and G25AP00379.
@@ A.1 Earthquake Catalog Generation
Data missingness, referred to in seismology as catalog (in)completeness, is the primary challenge faced with earthquake catalogs. It is an important and unavoidable feature, and is a result of how earthquakes are detected and characterised. Below, we briefly overview the process of generating an earthquake catalog to illustrate the data quality issues. In the subsequent section, we review catalog incompleteness and its potential impact on the performance and evaluation of forecasting models.
@@
Seismometers and Seismic Networks. A seismometer is an instrument that detects and records the vibrations caused by seismic waves (Stein and Wysession, 2009; Shearer, 2019). It consists of a sensor to detect ground motion and a recording system to log three-dimensional ground motion over time, typically vertical and horizontal velocities. Seismic networks, comprising multiple seismometers, monitor seismic activity at regional, national or global scales (see, e.g., (Woessner et al., 2010) and references therein). High-density networks with modern, sensitive equipment provide more detailed and accurate data, enhancing the ability to detect and analyse smaller and more distant earthquakes.
@@
From Waveforms to Phase Picking. The process of converting raw continuous seismic waveforms into useful earthquake data begins with phase picking, which identifies the arrival times of the primary (P) and secondary (S) waves of an earthquake. Historically, this was done manually, but now automated algorithms, such as the STA/LTA algorithm, detect wave arrivals by analyzing signal amplitude changes (Allen, 1982). Recent algorithms, such as machine learning classifiers (e.g. Zhu and Beroza, 2019; Lapins et al., 2021) and template-matching (e.g. Ross et al., 2019), can process much higher volumes of data efficiently and are often able to detect events of much smaller magnitudes.
@@
Earthquake Association and Location After phase picking, the next step is to associate phases from different seismometers with the same earthquake. Simple algorithms require at least four phase arrivals to be detected on different stations within a short time interval to declare an event. Once phases are associated, location estimation determines the earthquake’s hypocenter and origin time by minimizing travel-time residuals using linearized or global inversion algorithms (Thurber, 1985; Lomax et al., 2000). Given the potential for misidentified or mis-associated phase arrivals due to low signal-to-noise of small events or the near-simultaneous occurrence during very active aftershock sequences, an automated system typically first picks arrival times and determines a preliminary location, which is subsequently reviewed by a seismologist (e.g. Woessner et al., 2010, and references therein). Locations are typically reported as the geographical coordinates and depths where earthquakes first nucleated (hypocenters), although some catalogs report the centroid location, a central measure of the extended earthquake rupture.
@@
Earthquake Magnitude Calculation The magnitude of an earthquake quantifies the energy released at the source and was originally defined in the seminal paper by Richter (1935). The original definition, now referred to as the local magnitude (ML), is calculated from the logarithm of the amplitude of waves recorded by seismometers. This scale, however, "saturates" at higher magnitudes, meaning it underestimates magnitudes for various reasons. This led to introduction of the moment magnitude scale (Mw) (Hanks and Kanamori, 1979), which computes the magnitude based on the estimated seismic moment $M_{0}$, which can be related to the physical rupture process via
@@
where rigidity is a mechanical property of the rock along the fault, rupture area is the area of the fault that slipped, and slip is the distance the fault moved. Mw is determined seismologically via a spectral fitting process to the earthquake waveforms. In practice, it can be challenging to use a single magnitude scale for a broad range of magnitudes, therefore a range of scales may be present within a single catalog, and approximate magnitude conversion equations may be used to homogenize the scales (e.g. Herrmann and Marzocchi, 2021, and references therein).